OpenStudy (anonymous):

How do you simplify the square root of 180 and how do you know that it is in simplest form?

OpenStudy (anonymous):

to simplify square roots, separate them You know that \[\sqrt{4} * \sqrt{45} = \sqrt{180}\] square root of 4 is 2. so now you have \[2*\sqrt{45}\] Then you know that you can further simplify by separating more square roots \[2* \sqrt{9}*\sqrt{5}\] Square roots of 9 is 3. So you have \[2*3*\sqrt{5}\] \[= 6\sqrt{5}\]

OpenStudy (anonymous):

thank you study100

OpenStudy (anonymous):

welcome